OpenStudy (anonymous):

A boat traveled 64 miles downstream and back. The trip downstream took 4 hours. The trip back (upstream) took 8 hours. Find the speed of the boat n still water and the speed of the current. Please help me I don't remember how to solve problems like this. xo Thanks

OpenStudy (anonymous):

Alright, so I'm a bit fuzzy on how to do these sorts of problems, but I'll do my best to help! On the trip downstream, the boat traveled 64 miles in 4 hours, which is 16 miles per hour. On the trip upstream, the boat traveled 64 miles in 8 hours, which is 8 miles per hour. Let's say x is the speed of the boat in still water. Y is the speed of the current. Downstream: 16=x+y Upstream: 8=x-y Then set equal and solve!

OpenStudy (anonymous):

Thank you so so much! :)

OpenStudy (anonymous):

You're welcome!